OpenStudy (anonymous):

If the probability density of a random vribale is given by : f(x)= x for 0

OpenStudy (anonymous):

I already did the first one here: http://openstudy.com/users/jerjason#/updates/509d2b9fe4b0ac7e51942b75 I did this example in class and the answers my professor gave me for (a) was .3 for (a) and .5 for (b), my question is why am I only evaluating the integrals for f(x) = x and not f(x)=2-x as well?

OpenStudy (anonymous):

\[\int\limits_{.6}^{1.2}x dx = 1/2(x^2)]_{.6}^{1.2}=.54\] is what I got for (b) and rounded it is .5.

OpenStudy (sirm3d):

\[\large \int\limits_{0.6}^{1.2}f(x) dx=\int\limits_{0.6}^{1}f(x) dx+\int\limits_{1}^{1.2}f(x) dx\]

OpenStudy (sirm3d):

use \(\large f(x)=x\) for the first integral and \(\large f(x)=2-x\) for the second integral

OpenStudy (sirm3d):

as to your question of why only \(f(x)=x\) in (a), the function clearly states that \[\large f(x)=\begin{cases} x & 0<x<1,\\2-x & 1 \leq x< 2,\\0 & \text{elsewhere} \end{cases}\]

OpenStudy (sirm3d):

you must use \(f(x)=x\) when \(0 < x < 1\) since the interval given in (a) is a subset of this interval.

OpenStudy (anonymous):

Alright, I understand now. Thanks.